2. (4) Let W = span{(1,1,1), (-1,1,0)). Let v = (1,-1,2). Find the decomposition v = w; + W2, where we W and W, EW+.
Answer #1

Guren span { <1, 1, 1), (1, 2,0}} wt {ufu,vy * vew? =o Let u = (x,y,z) Ewt then, u in orthogonal to both {2,4A) and (-1,2,0)(+,-1,2) - (a-b, atb, a) + (1,9,-2x) (a-btn a-2x) ( 4,-1,2). arbte, 2 We get 1 a-ban 946+x= -4 ☆ a-2.1=2 -7-27 = 2 29 +22=0 a